magnetic susceptibility of copper is 0.5*10 calculate the magnetic moment per unit volume of copper,when it is subjected to a magnetic field of 10 G ?
1
Expert's answer
2013-07-22T10:11:28-0400
magnetic dipole moment per unit volume is
M=chi * H
where chi is magnetic susceptibility and H is magnetic field
Hence,
M = 0.5*10*10 = 50 G